# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v4

@@ Abstract.
Large Language Models (LLMs) have recently been leveraged for asset pricing tasks and stock trading applications, enabling AI agents to generate investment decisions from unstructured financial data. However, most evaluations of LLM timing-based investing strategies are conducted on narrow timeframes and limited stock universes, overstating effectiveness due to survivorship and data-snooping biases. We critically assess their generalizability and robustness by proposing FINSABER [^1] Data and code available at https://github.com/waylonli/FINSABER., a backtesting framework evaluating timing-based strategies across longer periods and a larger universe of symbols. Systematic backtests over two decades and 100+ symbols reveal that previously reported LLM advantages deteriorate significantly under broader cross-section and over a longer-term evaluation. Our market regime analysis further demonstrates that LLM strategies are overly conservative in bull markets, underperforming passive benchmarks, and overly aggressive in bear markets, incurring heavy losses. These findings highlight the need to develop LLM strategies that are able to prioritise trend detection and regime-aware risk controls over mere scaling of framework complexity.
@@
copyright: noneconference: ; Li et al.; 2025journalyear: 2025ccs: General and reference Evaluationccs: General and reference Empirical studiesccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Large language models (LLMs) are increasingly used in financial decision-making, especially for generating investment actions such as Buy, Hold, or Sell (ding2024largelanguagemodelagent; fatouros2025marketsenseai20enhancingstock). These so-called LLM timing-based investing strategies leverage LLMs’ ability to interpret historical and real-time data to autonomously trade. From sentiment-driven trading (zhang2024unveilingpotentialsentimentlarge) to sophisticated multi-agent systems (yu2023finmemperformanceenhancedllmtrading; finagentzhang2024multimodalfoundationagentfinancial), a growing body of work has explored the potential of LLMs as autonomous financial agents.
Backtesting is the standard method for assessing investment strategies, simulating them on historical data to evaluate profitability and robustness (chan2021quantitative). However, current LLM investing research suffers from fragmented, underdeveloped evaluation practices. Most studies assess performance over short periods, on few stock symbols, and often omit code release, limiting reproducibility. As summarised in Table 1, several recent methods evaluate over under a year, with fewer than ten stocks, and benchmark only against naïve baselines like Buy-and-Hold. Such short horizons and narrow stock universes lead to three well-documented sources of bias: survivorship bias (garcia1993survivorship), where delisted or failed stocks are omitted; look-ahead bias (chan2021quantitative), where future information inadvertently influences past decisions; and data-snooping bias (Bailey2015ThePO), where strategy performance is inflated through repeated testing on the same data. These biases can result in misleading performance assessments and undermine the validity of claimed improvements over traditional methods. This raises a central question: Can LLM-based investing strategies survive longer and broader robustness evaluations?
@@
While recent efforts such as wang2025quantbenchbenchmarkingaimethods and hu2025fintsbcomprehensivepracticalbenchmark have addressed benchmarking for deep learning (DL)-based trading and LLM-based time-series forecasting, comprehensive evaluation of LLM-based investing strategies remains unaddressed. Separately, FinBen (finben-2024) provides a thorough FinLLM benchmark covering multiple tasks, including decision-making. However, as a broad FinLLM benchmark, FinBen’s backtesting still relies on a limited, hand-picked symbol set, which contains the aforementioned biases and lacks a professional backtesting pipeline or systematic comparison with traditional strategies. To fill this gap, we introduce FINSABER, a comprehensive framework for benchmarking LLM timing-based investing strategies that supports longer backtesting periods, a broader and more diverse symbol universe, and explicit bias mitigation. Specifically, our main contributions are:
@@ 2. Related Works
Recent work using LLMs as investors directly employ LLMs to make investing decisions (ding2024largelanguagemodelagent). The most common approach leverages LLMs’ sentiment analysis capabilities, using either general-purpose LLMs (e.g., GPT, LLaMA (touvron2023llama), Qwen (bai2023qwentechnicalreport)) or fine-tuned financial variants like FinGPT (yang2023fingpt) to generate sentiment scores for trading decisions (lopezlira2023chatgptforecaststockprice; ruoxuportfolioperformance24; KIRTAC2024105227; zhang2024unveilingpotentialsentimentlarge). However, these approaches stop short of forming complete trading strategies, which require not only directional forecasts, but also realistic liquidity sizing for mitigating impact, development of execution rules for trade timing and risk management, and incorporation of trading costs.
More advanced approaches move beyond sentiment scores by summarising and reasoning over multi-source financial text. For example, fatouros2024largelanguagemodelsbeat introduce a memory module that stores summarised financial data, retrieved during trading to guide decisions. Similarly, LLMFactor (wang2024llmfactorextractingprofitablefactors) learns to extract profitable factors from historical news aligned with price movements and applies them to future market forecasts.
A growing body of work incorporates LLM-based agents (guo2024largelanguagemodelbased), where either one specialised agent or multiple collaborative agents are employed to perform financial analysis or predictions. Notable examples include FinMem (yu2023finmemperformanceenhancedllmtrading), FinAgent (finagentzhang2024multimodalfoundationagentfinancial), FinRobot (yang2024finrobotopensourceaiagent), TradExpert (ding2024tradexpertrevolutionizingtradingmixture), FinCon (yu2024fincon), TradingAgents (xiao2024tradingagents) and MarketSenseAI 2.0 (fatouros2025marketsenseai20enhancingstock). Some models also incorporate reinforcement learning (RL) for iterative self-improvement (RePEc:arx:papers:2310.05627; 10.1145/3589334.3645611).
@@ 4. Biases and Robustness Challenges in Backtesting LLM Investors
Robust evaluation of financial strategies demands carefully designed backtests. Unlike typical machine learning tasks with large, clean datasets, financial data is noisy, nonstationary, and limited in scope. As a result, backtests are especially prone to three major sources of bias: survivorship bias, look-ahead bias, and data-snooping bias, each of which can inflate perceived performance and lead to misleading conclusions (chan2021quantitative).
@@
This occurs when backtests include only currently active stocks while ignoring delisted or bankrupt assets. Such omissions systematically overstate returns and understate risk (joubert2024three). A common cause is using today’s S&P 500 constituents as the historical investment universe. This practice introduces what garcia1993survivorship call “preinclusion bias”, also a form of look-ahead bias where future index membership influences past decisions. The impact is well-documented: grinblatt1989mutual and elton1996survivor estimate annual return distortions between 0.1% and 0.9%, and brown1992survivorship show that even small distortions can misrepresent performance persistence.
@@
Look-ahead bias arises when a strategy uses information that would not have been known at the time of decision-making (chan2021quantitative). This includes selecting features, parameters, or symbols based on full-period outcomes, thereby introducing future knowledge into the backtest.
@@
Also known as multiple testing bias, this occurs when repeated experimentation on the same dataset leads to overfitting. In finance, where sample sizes are small and the signal-to-noise ratio is very low, this bias is particularly problematic. Bailey2015ThePO showed that evaluating strategies on overlapping data inflates false positive rates, and that standard hold-out validation techniques often fail to guard against this issue.
@@
Addressing these biases requires evaluating strategies across longer periods and broader asset universes. For daily trading, at least three years of data is generally recommended, while weekly and monthly strategies benefit from 10 to 20 years or more (Bailey2015ThePO). 10.1093/rfs/hhj020 tested pairs trading on 40 years of daily data, but do2010does extended this to 48 years and found profitability declined, highlighting the need for long-term evaluation. Likewise, recent deep learning models in finance rely on multi-year datasets to ensure robustness (fengtemporal; ijcai2022p0551; CI-STHPAN).
@@ 5. FINSABER
Multi-source Data for LLM Benchmarking.
@@
We incorporate a comprehensive collection of strategies across multiple paradigms to ensure robust benchmarking. The timing-based strategies include open-source LLM investors (FinMem (yu2023finmemperformanceenhancedllmtrading), FinAgent (finagentzhang2024multimodalfoundationagentfinancial)), traditional rule-based approaches (Buy and Hold, Moving Average Crossover, Bollinger Bands (bollinger2002bollinger), Trend Following (Wilcox2009trendfollowing)), ML/DL forecaster-based methods (ARIMA, XGBoost), and RL-based strategies (A2C, PPO, TD3, SAC implemented via FinRL (finrl-2022) framework). Selection-based strategies encompass random K selection, Momentum Factor Selection (based on past returns), Volatility Effect Selection (selecting low-volatility stocks), and the stocks selection agent from the FinCon (yu2024fincon) framework. This diverse strategy base enables comprehensive performance comparison across different methodological approaches while maintaining extensibility for custom implementations. More technical details of the strategies are available in Appendix B.
@@
FINSABER adopts three categories of evaluation metrics: return, risk, and risk-adjusted performance. Return metrics measure profitability and include Annualised Return (AR) and Cumulative Return (CR). Risk metrics quantify uncertainty and downside exposure, including Annualised Volatility (AV) and Maximum Drawdown (MDD). Risk-adjusted metrics assess capital efficiency and include the Sharpe Ratio (SPR) and Sortino Ratio (STR).
High returns alone do not imply strategy quality. Risk-adjusted metrics such as SPR and STR are more informative, especially in finance where capital efficiency and downside risk are critical (chan2021quantitative). These metrics are standard in the literature (cont2001empirical; demiguel2009optimal) and widely used in recent LLM-based investing benchmarks (finagentzhang2024multimodalfoundationagentfinancial; yu2024fincon). Formal definitions and formulas are provided in Appendix C.
@@ 6.1. Pitfalls of Selective Evaluation
Rule
Based
Buy and Hold -0.342 -20.483 -52.729 55.910 1.326 43.079 -20.184 41.523 -0.460 -13.250 -31.546 35.624 0.974 21.171 -14.192 28.327
@@
Rule
Based
Buy and Hold 0.630 0.915 37.767 -50.839 45.243 0.622 0.952 23.919 -48.119 41.703
@@
Rule
Based
Buy and Hold 0.551 0.829 15.997 -36.842 30.860 0.461 0.620 11.238 -25.463 21.791
@@
6.2. Fair Comparisons with the Composite Approach
To overcome the aforementioned biases, we introduce the Composite evaluation setup within FINSABER. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use four unbiased stock selection approaches from the strategies base (details in Appendix B): Random Five, Momentum Factor (Muller31032010), Volatility Effect (volatility-effect), and the FinCon Selection Agent in the FinCon (yu2024fincon) framework.
@@ 6.2. Fair Comparisons with the Composite Approach
To mitigate survivorship bias, we use historical constituent lists, specifically S&P 500 for US market, at each evaluation period’s start and explicitly include delisted symbols. To address data-snooping bias, we evaluate a large and diversified symbol universe: 91, 84, and 63 total distinct symbols for Random Five, Momentum-based, and Volatility-based selection, respectively. These counts reflect all unique symbols encountered across rolling windows, where stocks are reselected in each window, preventing cherry-picking and short-horizon bias.
@@
Rule
Based
Buy and Hold 0.315 0.456 6.694 -35.130 27.410 0.384 0.694 9.916 -32.596 37.421
@@
Rule
Based
Buy and Hold 0.703 1.291 7.898 -14.146 14.720 0.389 0.671 6.940 -30.943 41.710
@@
Table 4 summarises these comprehensive evaluations. Results obtained through this unbiased and systematic approach further validate our previous findings from the selected-four evaluation. Specifically, both the Random Five and Momentum-based selections reinforce the conclusion that the previously claimed superiority of LLM investors is largely driven by selective evaluation setups. For instance, in the Random Five setup, Buy and Hold, ATR Band and ARIMA outperform FinMem and FinAgent in terms of risk-adjusted metrics. Similarly, ARIMA and simple rule-based strategies often perform better than LLM-based methods under the Momentum-based selection. In the Volatility-based selection, traditional methods dominate even more clearly: Buy and Hold achieves the highest Sharpe (0.703), Sortino (1.291), and AR (7.898%), while PPO and ARIMA again show strong all-round performance. LLM-based methods lag behind, with FinAgent offering moderate returns but lower Sharpe (0.241) and larger drawdowns. Notably, our reported LLM performances do not adjust for potential data leakage: given the use of pretrained models like GPT-4o, the LLMs may have seen parts of the data during training—yet they still fail to outperform traditional strategies under fair evaluation, casting further doubt on their real-world advantage.

### v5

@@ Abstract.
Large Language Models (LLMs) have recently been leveraged for asset pricing and stock trading applications, enabling AI agents to generate investment decisions from unstructured financial data. However, most evaluations of LLM timing-based investing strategies are conducted on narrow timeframes and limited stock universes, overstating effectiveness due to survivorship and data-snooping biases. We critically assess their generalisability and robustness by proposing FINSABER [^1] Data and code available at https://github.com/waylonli/FINSABER., a backtesting framework evaluating timing-based strategies across longer periods and a larger universe of symbols. Systematic backtests over two decades and 100+ symbols reveal that previously reported LLM advantages deteriorate significantly under broader cross-section and over a longer-term evaluation. Our market regime analysis further demonstrates that LLM strategies are overly conservative in bull markets, underperforming passive benchmarks, and overly aggressive in bear markets, incurring heavy losses. These findings highlight the need to develop LLM strategies that are able to prioritise trend detection and regime-aware risk controls over mere scaling of framework complexity.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3785702isbn: 979-8-4007-2258-5/2026/08ccs: General and reference Evaluationccs: General and reference Empirical studiesccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Large language models (LLMs) are increasingly used in financial decision-making, especially for generating investment actions such as Buy, Hold, or Sell (Ding et al., 2024a; Fatouros et al., 2025). These so-called LLM timing-based investing strategies leverage LLMs’ ability to interpret historical and real-time data to autonomously trade. From sentiment-driven trading (Zhang et al., 2023) to sophisticated multi-agent systems (Yu et al., 2023; Zhang et al., 2024), a growing body of work has explored the potential of LLMs as autonomous financial agents.
Backtesting is the standard method for assessing investment strategies, simulating them on historical data to evaluate profitability and robustness (Chan, 2021; Li and Ma, 2025). However, current LLM investing research suffers from fragmented, underdeveloped evaluation practices. Most studies assess performance over short periods, on few stock symbols, and often omit code release, limiting reproducibility. As summarised in Table 1, several recent methods evaluate over under a year, with fewer than ten stocks, and benchmark only against naïve baselines like Buy-and-Hold. Such short horizons and narrow stock universes lead to three sources of bias: survivorship bias (Garcia and Gould, 1993), where delisted or failed stocks are omitted; look-ahead bias (Chan, 2021), where future information inadvertently influences past decisions; and data-snooping bias (Bailey et al., 2015), where strategy performance is inflated through repeated testing on the same data. These biases can result in misleading performance assessments and undermine the validity of claimed improvements over traditional methods. This raises a central question: Can LLM-based investing strategies survive longer and broader robustness evaluations?
@@
While recent efforts such as Wang et al. (2025) and Hu et al. (2025) have addressed benchmarking for deep learning (DL)-based trading and LLM-based time-series forecasting, comprehensive evaluation of LLM-based investing strategies remains unaddressed. Separately, FinBen (Xie et al., 2024) provides a thorough FinLLM benchmark covering multiple tasks, including decision-making. However, as a broad FinLLM benchmark, FinBen’s backtesting still relies on a limited, hand-picked symbol set, which contains the aforementioned biases and lacks a professional backtesting pipeline or systematic comparison with traditional strategies. To fill this gap, we introduce FINSABER, a comprehensive framework for benchmarking LLM timing-based investing strategies that supports longer backtesting periods, a broader and more diverse symbol universe, and explicit bias mitigation. Specifically, our main contributions are:
@@ 2. Related Works
Recent work using LLMs as investors directly employ LLMs to make investing decisions (Ding et al., 2024a). The most common approach leverages LLMs’ sentiment analysis capabilities, using either general-purpose LLMs (e.g., GPT, LLaMA, Qwen) or fine-tuned financial variants like FinGPT (Yang et al., 2023) to generate sentiment scores for trading decisions (A. Lopez-Lira and Y. Tang (2023); R. Wu (2024); 1; H. Zhang, F. Hua, C. Xu, H. Kong, R. Zuo, and J. Guo (2023)). However, these approaches stop short of forming complete trading strategies, which require not only directional forecasts, but also realistic liquidity sizing for mitigating impact, development of execution rules for trade timing and risk management, and incorporation of trading costs.
More advanced approaches move beyond sentiment scores by summarising and reasoning over multi-source financial text. For example, Fatouros et al. (2024) introduce a memory module that stores summarised financial data, retrieved during trading to guide decisions. Similarly, LLMFactor (Wang et al., 2024) learns to extract profitable factors from historical news aligned with price movements and applies them to future market forecasts.
A growing body of work incorporates LLM-based agents (Guo et al., 2024), where either one specialised agent or multiple collaborative agents are employed to perform financial analysis or predictions. Notable examples include FinMem (Yu et al., 2023), FinAgent (Zhang et al., 2024), FinRobot (Yang et al., 2024), TradExpert (Ding et al., 2024b), FinCon (Yu et al., 2024), TradingAgents (Xiao et al., 2024) and MarketSenseAI 2.0 (Fatouros et al., 2025). Some models also incorporate reinforcement learning (RL) for iterative self-improvement (Ding et al., 2023; Koa et al., 2024).
@@ 4. Biases and Robustness Challenges in Backtesting LLM Investors
Robust evaluation of financial strategies demands carefully designed backtests. Unlike typical machine learning tasks with large, clean datasets, financial data is noisy, nonstationary, and limited in scope. As a result, backtests are especially prone to three major sources of bias: survivorship bias, look-ahead bias, and data-snooping bias, each of which can inflate perceived performance and lead to misleading conclusions (Chan, 2021).
@@
This occurs when backtests include only currently active stocks while ignoring delisted or bankrupt assets. Such omissions systematically overstate returns and understate risk (Joubert et al., 2024). A common cause is using today’s S&P 500 constituents as the historical investment universe. This practice introduces what Garcia and Gould (1993) call “preinclusion bias”, also a form of look-ahead bias where future index membership influences past decisions. The impact is well-documented: Grinblatt and Titman (1989) and Elton et al. (1996) estimate annual return distortions between 0.1% and 0.9%, and Brown et al. (1992) show that even small distortions can misrepresent performance persistence.
@@
Look-ahead bias arises when a strategy uses information that would not have been known at the time of decision-making (Chan, 2021). This includes selecting features, parameters, or symbols based on full-period outcomes, thereby introducing future knowledge into the backtest.
@@
Also known as multiple testing bias, this occurs when repeated experimentation on the same dataset leads to overfitting. In finance, where sample sizes are small and the signal-to-noise ratio is very low, this bias is particularly problematic. Bailey et al. (2015) showed that evaluating strategies on overlapping data inflates false positive rates, and that standard hold-out validation techniques often fail to guard against this issue.
@@
Addressing these biases requires evaluating strategies across longer periods and broader asset universes. For daily trading, at least 3 years of data is generally recommended, while weekly and monthly strategies benefit from 10 to 20 years or more (Bailey et al., 2015). Gatev et al. (2006) tested pairs trading on 40 years of daily data, but Do and Faff (2010) extended this to 48 years and found profitability declined, highlighting the need for long-term evaluation. Likewise, recent deep learning models in finance rely on multi-year datasets to ensure robustness (Feng et al., 2019; Wang et al., 2022).
@@ 5. FINSABER
Multi-source Data.
@@
We incorporate a comprehensive collection of strategies across multiple paradigms to ensure robust benchmarking. The timing-based strategies include open-source LLM investors (FinMem (Yu et al., 2023), FinAgent (Zhang et al., 2024)), traditional rule-based approaches (Buy and Hold, Moving Average Crossover, Bollinger Bands (Bollinger, 2002), Trend Following (Wilcox et al., 2005)), ML/DL forecaster-based methods (ARIMA, XGBoost), and RL-based strategies (A2C, PPO, TD3, SAC implemented via FinRL (Liu et al., 2022) framework). Selection-based strategies encompass random K selection, Momentum Factor Selection (based on past returns), Volatility Effect Selection (selecting low-volatility stocks), and the stocks selection agent from the FinCon (Yu et al., 2024) framework. This diverse strategy base enables comprehensive performance comparison across different methodological approaches while maintaining extensibility for custom implementations. More technical details of the strategies are available in Appendix B.
@@
FINSABER adopts three categories of evaluation metrics: return, risk, and risk-adjusted performance. Return metrics measure profitability, including Annualised Return (AR) and Cumulative Return (CR). Risk metrics quantify uncertainty and downside exposure, including Annualised Volatility (AV) and Maximum Drawdown (MDD). Risk-adjusted metrics assess capital efficiency, including the Sharpe Ratio (SPR) and Sortino Ratio (STR).
High returns alone do not imply strategy quality. Risk-adjusted metrics such as SPR and STR are more informative, especially in finance where capital efficiency and downside risk are critical (Chan, 2021). These metrics are standard in the literature (Cont, 2001; DeMiguel et al., 2007) and widely used in recent LLM-based investing benchmarks (Zhang et al., 2024; Yu et al., 2024). Formal definitions and formulas are provided in Appendix C.
@@ 6.1. Pitfalls of Selective Evaluation
Rule Based Buy and Hold -0.342 -20.483 -52.729 55.910 1.326 43.079 -20.184 41.523 -0.460 -13.250 -31.546 35.624 0.974 21.171 -14.192 28.327
@@
Rule Based Buy and Hold 0.630 0.915 37.767 -50.839 45.243 0.622 0.952 23.919 -48.119 41.703
@@
Rule Based Buy and Hold 0.551 0.829 15.997 -36.842 30.860 0.461 0.620 11.238 -25.463 21.791
@@
6.2. Fair Comparisons with Composite Approach
To overcome the aforementioned biases, we introduce the Composite evaluation setup within FINSABER. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use four unbiased stock selection approaches from the strategies base (details in Appendix B): Random Five, Momentum Factor (Muller and Ward, 2010), Volatility Effect (Blitz and Vliet, 2007), and the FinCon Selection Agent in the FinCon (Yu et al., 2024) framework.
@@ 6.2. Fair Comparisons with Composite Approach
To mitigate survivorship bias, we use historical constituent lists, specifically S&P 500 for US market, at each evaluation period’s start and explicitly include delisted symbols. To address data-snooping bias, we evaluate a large and diversified symbol universe: 91, 84, 63, and 80 total distinct symbols for Random Five, Momentum-based, Volatility-based selection, and FinCon Selection Agent respectively. These counts reflect all unique symbols encountered across rolling windows, where stocks are reselected in each window, preventing cherry-picking and short-horizon bias.
@@
Rule Based Buy and Hold 0.315 0.456 6.694 -35.130 27.410 0.384 0.694 9.916 -32.596 37.421
@@
Rule Based Buy and Hold 0.703 1.291 7.898 -14.146 14.720 0.389 0.671 6.940 -30.943 41.710
@@
Table 4 summarises these comprehensive evaluations. Results obtained through this unbiased and systematic approach further validate our previous findings from the selected-four evaluation. Specifically, both the Random Five and Momentum-based selections reinforce the conclusion that the previously claimed superiority of LLM investors is largely driven by selective evaluation setups. For instance, in the Random Five setup, Buy and Hold, ATR Band and ARIMA outperform FinMem and FinAgent in terms of risk-adjusted metrics. Similarly, ARIMA and simple rule-based strategies often perform better than LLM-based methods under the Momentum-based selection. In the Volatility-based selection, traditional methods dominate even more clearly: Buy and Hold achieves the highest Sharpe (0.703), Sortino (1.291), and AR (7.898%), while PPO and ARIMA again show strong all-round performance. LLM-based methods lag behind, with FinAgent offering moderate returns but lower Sharpe (0.241) and larger drawdowns. Notably, our reported LLM performances do not adjust for potential data leakage: given the use of pretrained models like GPT-4o, the LLMs may have seen parts of the data during training, but they still fail to outperform traditional strategies under fair evaluation, casting further doubt on their real-world advantage.